OpenStudy (anonymous):

can someone please help me with 4 questions ?

OpenStudy (anonymous):

@superman38

OpenStudy (anonymous):

shure

OpenStudy (anonymous):

which fraction is equivalent to 56% 14/25 7/8 5/6 56/10

OpenStudy (anonymous):

its either a or d

OpenStudy (anonymous):

ok do u know for sure which it is ..can u work on it please ill be back just check it over ig and ill be back

OpenStudy (anonymous):

ight

OpenStudy (anonymous):

@baconbaby1334

OpenStudy (anonymous):

@lil_chukie

OpenStudy (anonymous):

it is A

OpenStudy (anonymous):

how did u get the answer though can u explain a little of how u would get the answer? @chethus

OpenStudy (anonymous):

yea a

OpenStudy (anonymous):

yeah sure...to get any fraction in percentage we simply multiply that fraction by hundred...here \[(14\div25)\times100\]

OpenStudy (anonymous):

that is 14 x 4=56...so that is the rite option

OpenStudy (anonymous):

i know u still have some doubt...

OpenStudy (anonymous):

oh ok thx so much can yall help me with a few more maybe @superman38 @chethus @lil_chukie

OpenStudy (anonymous):

yup....but pls make it fast sis...:)

OpenStudy (anonymous):

What is 5% of 78 0.39 3.9 39 390

OpenStudy (anonymous):

can u help superman38?...@chethus

OpenStudy (lynfran):

3.9

OpenStudy (anonymous):

ok can u help with 2 more maybe?

OpenStudy (lynfran):

sure

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

12% of what number is 180 1500 15 21.6 150 when you do percentage do u add subtract divide? I was thinking D ?

OpenStudy (lynfran):

1500

OpenStudy (lynfran):

12/100*1500=180

OpenStudy (anonymous):

so A

OpenStudy (lynfran):

or u can also approach it as 100/12*180=1500

OpenStudy (lynfran):

yes A

OpenStudy (anonymous):

The regular price of a music player is $59.50. It was on sale for 24% off How much was the discount I have no answers ...u have to figure it out $ ?

OpenStudy (lynfran):

the discount 24/100*59.50=$14.28 and the new price would be 59.50-14.28=$45.22 if ur curious to know this part..

OpenStudy (anonymous):

ok thx one more

OpenStudy (lynfran):

ok

OpenStudy (anonymous):

Miriam bought a digital camera that cost $140.The sales tax n her state is 7%. what I the final amount that Miriam paied for the camera?

OpenStudy (anonymous):

Miriam bought a digital camera that cost $140.The sales tax n her state is 7%. what I the final amount that Miriam paied for the camera?

OpenStudy (lynfran):

7/100*140=9.8 140+9.8=149.8 the answer $149.8 she will pay for it

OpenStudy (anonymous):

ok thx so much can I fan u?

OpenStudy (lynfran):

sure